# Euclid: A blue galaxy population and a brightest cluster galaxy in the making in a MaDCoWS2 galaxy cluster candidate††thanks: This paper is published on behalf of the Euclid Consortium

### 07 Mar 2026 v2

@@ 2.2.2 MER images
The MER catalogue pipeline performs well for most galaxies, but tends to overcorrect or suppress low-surface brightness features like tidal tails (Euclid Collaboration: Aussel et al., 2025; Euclid Collaboration: Roster et al., 2025). Furthermore, the single Sérsic and circular aperture photometry currently available are poorly suited to irregular galaxies. Thus, rather than rely on the MER catalogue photometry of the BCG complex, we downloaded Euclid and DECam level 2 mosaics of the Puddle cluster, and trimmed them to a more manageable size of . We estimated the flux of the BCG with sep (Barbary, 2018), using a custom aperture (see Sect. 3.3). The level 2 DECam mosaics in the Euclid Science Archive System are already resampled to the pixel scale of the Euclid images.
@@ 2.3 Spitzer data
The Spitzer Space Telescope (Werner et al., 2004) was a near- and mid-infrared observatory launched in August 2003 with three instruments onboard: the Infrared Array Camera (IRAC, Fazio et al., 2004), the Multiband Infrared Photometer for Spitzer (MIPS; Rieke et al., 2004), and the Infrared Spectrograph (IRS, Houck et al., 2004). IRAC had four bands, usually called ‘channels,’ centred on , , , and ; MIPS had three bands, centred on , , and . However, after Spitzer exhausted its liquid helium reserve in May 2009, only the two bluest bands of the IRAC camera could be used.
The Cosmic Dawn Survey (Euclid Collaboration: Moneti et al., 2022, see also Euclid Collaboration: Zalesky et al. 2025) is an IRAC survey covering the three Euclid Deep Fields and several calibration fields. Its image mosaics are based on both dedicated and archival data in the four IRAC channels, and correspond to about 11% of the telescope total mission time (Euclid Collaboration: Moneti et al., 2022). A source catalogue is available; however, the aperture sizes are inappropriate for the shape of the BCG. Furthermore, the BCG complex is blended with many neighbouring galaxies at and and is barely detected at and . We thus estimated the flux directly from the mosaics (see Sect. 3.3).
The coordinates of the brightest core in the BCG complex correspond ( separation) to a point source detected at in the Spitzer Wide-area InfraRed Extragalactic Survey (SWIRE; see Lonsdale et al., 2003, 2004). Given the relative isolation of this source (the closest MIPS source is away) and the FHWM of the MIPS point-spread function (PSF; Rieke et al., 2004), we adopt the SWIRE flux here.
@@ 2.4 Keck MOSFIRE spectroscopy
The Euclid Q1 release contains NISP slitless multiobject spectra taken with the red grism (Euclid Collaboration: Copin et al., 2025; Euclid Collaboration: Jahnke et al., 2025). However, the reduction of the grism data was not optimal in Q1; most of the identified issues are being addressed for the Euclid DR1. In the Puddle cluster field of view, an average of 55.6% of the pixels of the calibrated unidimensional spectra (excluding the edges) are flagged for suspicious behaviour.
@@
Spectroscopic data were reduced with the PypeIt data reduction pipeline version 1.17.1 (Prochaska et al., 2020), following the standard calibration and extraction process. The reduction process included bias subtraction, flat-field correction, and wavelength calibration employing arc lamp frames. Sky subtraction was performed using the PypeIt algorithm, which models the sky background on a slit-by-slit basis to account for variations across the field. PypeIt outputs reduced and co-added 2D spectra, from which we extract 1D spectra using a constant extraction width as a function of wavelength with an object-finding algorithm. The extraction width is calculated based on a PypeIt-based optimal aperture mask maximising the S/N. We examined the 1D and 2D spectra to locate emission lines.
@@ 3 Analysis
The detection of the Puddle cluster in MaDCoWS2 and Euclid (Euclid Collaboration: Bhargava et al., 2025) indicates the presence of a sizeable overdensity. MOSFIRE spectroscopy provides spectroscopic redshifts centred at for the BCG (see Sect. 3.2) and a galaxy at away from the BCG complex, which may be part of the cluster infalling region. In this section, we explore the candidate cluster properties with the photometric and spectroscopic data sets. When necessary, we assume that the redshifts of the cluster and the BCG are .
@@ 3.1 The galaxy population
Figure 2 presents the colour-magnitude diagram (CMD) of the Puddle cluster. The galaxies out to away (about Mpc) from the proto-BCG are shown as green points, and those between and are in black. The red sequence colour is 2.44 at (pink line), assuming Solar metallicity. It is computed with Bagpipes (Carnall et al., 2018) with a model consisting of a single burst of star formation at followed by passive evolution. Combining this model with the Mancone et al. (2010) luminosity function for the IRAC channel, we compute a characteristic magnitude () of 21.4 in band. A formation redshift of 3 generates similar results: a red sequence at and .
The central structure is detected as four galaxies by the MER pipeline. However, given the irregular shape and the presence of substantial diffuse emission, the MER photometry is unreliable for these four sources. Thus, we remove those four detections from our CMDs and perform our own measurement of the proto-BCG flux in Sect. 3.3. The cyan dot in Fig. 2 represents the position of the central structure in the CMD.
The red sequence of our CMD appears sparsely populated compared with clusters confirmed at (e.g. Papovich et al., 2010; Andreon et al., 2014; Webb et al., 2015b; Strazzullo et al., 2016; Willis et al., 2020). We see in particular a lack of red galaxies at , which is unusual among the known high-redshift clusters: massive galaxies are more susceptible to internal quenching (e.g. Peng et al., 2010b) and thus tend to be redder than less massive ones even at (e.g. Kawinwanichakij et al., 2017).
@@ 3.1.1 Background subtraction and completeness correction
To assess the importance of the galaxy overdensity associated with the Puddle cluster and its location in the CMD diagram, we estimate the contribution of field galaxies and subtract it from our CMD. To measure this ‘background,’ we randomly selected 100 non-overlapping circular regions () in the Fornax field. We compute the average density of galaxies in these fields as a function of magnitude and colour, using only the galaxies that are at least -detected in both and bands. For each magnitude and colour, we then subtract this average density from the galaxy density in the central region of the Puddle cluster. The background subtraction uncertainties are determined from the standard deviation of the galaxy densities in the random fields.
@@
We also use the 100 random fields described above to determine completeness corrections. Our computation follow the same principle than the completeness correction described in the Appendix A.1. of van der Burg et al. (2020). Figure 3 presents a visual summary of the process. We fit an exponential to the highly complete part of the distribution in each band empirically determined as and . It should be noted that since the -band completeness is the most critical one, we have been especially conservative in our selection of the regime which we consider highly complete. We then compute a completeness factor for magnitudes greater than 23 () or 23.5 () by taking the ratio of the observed galaxy count to the count predicted by the best fit. We assume a completeness factor of one for and . The total 50% completeness, shown on Fig. 2 as a purple line, is estimated by multiplying the and -band completeness factor for a given set of and magnitudes. This multiplication relies on the assumption that there is no correlation between detections in the and bands. Since that may not be true, we limit the rest of our analysis to , a regime where the completeness correction does not fall below 34% and is entirely determined by the -band completeness (i.e. the completeness correction is entirely determined by colour). The completeness uncertainties are propagated from the uncertainties on the best fit parameters. For example, the correction for a -band completeness of 0.5 is .
@@
Figure 4 presents the impact of these corrections on the galaxy population of the Puddle cluster candidate. The bottom left panel shows the background-subtracted, completeness-corrected distribution. The 50% -band completeness limits is shown as a cyan line. To avoid large (and uncertain) completeness corrections, we limit the CMD in the bottom left panel to ( is equivalent to on the red sequence and to a 76% complete band). We limit the colour to , as the background-subtracted galaxy count is consistent with zero for redder colours. The other panels show the galaxy count before and after the application of the completeness correction as a function of (top left) and colour (bottom right). For , the excess count within of the BCG is galaxies above the field level. It increases to galaxies after the application of the completeness correction. The red sequence location is indicated on the top left and bottom panels. The count uncertainties are based on the propagation of the density subtraction and the completeness uncertainties when relevant.
@@
As noted above, there are galaxies ( after the completeness correction) in excess compared to the field level in the red sequence and very few luminous red galaxies outside of the BCG complex ( galaxies with ). The bulk of the overdensity is about 1.5 magnitudes bluer than the red sequence population, and fainter than . With Bagpipes, we compute that corresponds to a Myr old stellar population, assuming Solar metallicity and no attenuation. While this blue colour suggests widespread star formation, we do not attempt to determine the cluster specific SFR. The lack of bright red galaxies and the imprecise photometric redshifts of Euclid at , are probably the two main factors explaining the difference between our galaxy counts and the Euclid Collaboration: Bhargava et al. (2025) richness estimate.
@@ 3.2 Spectroscopy
We detect continuum emission of varying robustness in 14 of the 18 MOSFIRE slitlets. We also identify unambiguous isolated emission lines, most likely H , in three slitlets: one slitlet targeted the BCG complex, one slitlet targeted a source approximately from the MaDCoWS2 centroid, and another slitlet far from the MaDCoWS2 centroid has two serendipitous emission lines. In addition, faint, tentative emission lines are detected in three other slitlets. Based on the latter, we estimate a line detection threshold of at the central wavelengths of the -band grism, corresponding to a star-formation rate of assuming a standard conversion factor between star-formation rate and H line luminosity (Hao et al., 2011). At m, the wavelength of the emission line detected from the BCG complex, the atmosphere absorbs of the flux, implying a rough star-formation rate detection threshold of 8. The presence of faint traces and tentative emission lines suggests that exposure time calculation were perhaps too optimistic. We also note that the outcome of our observation is comparable to Stanford et al. (2012) attempt to confirm IDCS J1426.5+3508, a cluster, with spectra from the Low-Resolution Imaging Spectrograph on Keck I. On 16 targets, they confirmed two members. However, the main challenge in our data reduction is the subtraction of the strong near-infrared sky lines (e.g. Webb et al., 2015b; Balogh et al., 2017; Belli et al., 2019), while Stanford et al. (2012) main limitation is the faintness of quiescent galaxies in the optical (which corresponds to rest-frame ultraviolet at ).
@@
The emission region in Fig. 5 is broad, spanning roughly Å and presents a complex structure. Such a broad range suggests three possibilities: a single H line from an AGN; at least two sources of H emission along the same line-of-sight; or H emission blended with the [N ii] doublet. We examine these possibilities by modelling the spectrum as one, two, and three Gaussians, weighting the contribution of each pixel by the inverse of its variance. We compute the reduced and Akaike Information Criterion (AIC, e.g. Akaike, 1974; Liddle, 2007) of each fit. The results are presented in Table 1. The middle and lowest panels of Fig. 5 show the one-, two-, and three-Gaussian decompositions in blue, pink, and purple respectively. Due to the limited S/N of the spectrum, we do not attempt to subdivide the H emission into broad and narrow components.
The simplest case, the single Gaussian model, yields the highest reduced and AIC. Both measures strongly suggest that the two- or three-Gaussians models are better representations of the spectrum. The components of the three-Gaussian model are not treated independently: we fix the amplitude ratio of the [N ii] doublet as , in agreement with theoretical (e.g. Galavís et al., 1997; Storey and Zeippen, 2000; Bon et al., 2025) and observational (Dočinović et al., 2023) determinations, and force its members to have the same width. We force all three emission lines to have the same redshifts. We also test a fit in which we relax the last requirement, allowing the H emission to lie at a different redshift than the [N ii] doublet. We find very similar redshifts () and the FWHM of each line is not significantly changed compared to the previous fit. We discuss the interpretation of the various fits in Sect. 4.2.1.
@@ 3.3 The BCG complex
The right panel of Fig. 1 presents a tricolour image (, , and ) of the BCG complex of the Puddle cluster, with the MOSFIRE slit and the custom aperture used for the photometry overlaid (see Sect. 3.3.1). This system has a complex morphology, with diffuse emission embedding at least six cores and extending in a kind of tail toward the southwest. There is a seventh core within the tail emission. The projected distance between cores is kpc, and in the case of the six central cores kpc. Along the axis of the diffuse tail, the system is kpc across.
@@ 3.3.1 Euclid and DECam photometry
To estimate the magnitude of the BCG complex, our first step is to determine a suitable aperture to measure the system photometry. To do so, we use sep.extract to compute the segmentation map in the NISP band, setting our detection threshold at with a minimum area of 30 pixels. We then assign a value of one to the pixel corresponding to the BCG complex detections and remove all other detections in our segmentation map. The resulting detection is very irregularly shaped, so we use a Gaussian kernel with 3 pixels to smooth the detection shape, retaining only the pixels with a value of 0.25 or more. We then use the resulting shape (shown as white contours in the right panel of Fig. 1) to measure the photometry in all Euclid and DECam bands.
To estimate the photometric uncertainties, we adopt a statistical approach: we find all detections in the image of interest, with a detection threshold at with a minimum area of 5 pixels. We then randomly displace the footprint of the central merger until we find a position with no overlapping detections. We measure the sky flux in that position and repeat the process for 100 random empty sky patches. We take the standard deviation of the sky fluxes as our sky uncertainties, which we then add in quadrature to the Poisson noise of the aperture photometry.
We do not compute aperture corrections for DECam: as long as the defined aperture contains most of the flux in every band, the aperture corrections should be negligible. This is the case for all DECam bands. In 3 segmentation maps, the BCG complex, if detected, is entirely enclosed by our custom aperture.

### 13 Mar 2026 v3

@@ 2.2.2 MER images
The MER catalogue pipeline performs well for most galaxies, but tends to overcorrect or suppress low-surface brightness features like tidal tails (Euclid Collaboration: Aussel et al. 2025; Euclid Collaboration: Roster et al. 2025). Furthermore, the single Sérsic and circular aperture photometry currently available are poorly suited to irregular galaxies. Thus, rather than rely on the MER catalogue photometry of the BCG complex, we downloaded Euclid and DECam level 2 mosaics of the Puddle cluster, and trimmed them to a more manageable size of . We estimated the flux of the BCG with sep (Barbary 2018), using a custom aperture (see Sect. 3.3). The level 2 DECam mosaics in the Euclid Science Archive System are already resampled to the pixel scale of the Euclid images.
@@ 2.3 Spitzer data
The Spitzer Space Telescope (Werner et al. 2004) was a near- and mid-infrared observatory launched in August 2003 with three instruments on board: the Infrared Array Camera (IRAC, Fazio et al. 2004), the Multiband Infrared Photometer for Spitzer (MIPS; Rieke et al. 2004), and the Infrared Spectrograph (IRS, Houck et al. 2004). IRAC had four bands, usually called ‘channels’, centred on , , , and ; MIPS had three bands, centred on , , and . However, after Spitzer exhausted its liquid helium reserve in May 2009, only the two bluest bands of the IRAC camera could be used.
The Cosmic Dawn Survey (Euclid Collaboration: Moneti et al. 2022, see also Euclid Collaboration: Zalesky et al. 2025) is an IRAC survey covering the three Euclid Deep Fields and several calibration fields. Its image mosaics are based on both dedicated and archival data in the four IRAC channels, and correspond to about 11% of the telescope total mission time (Euclid Collaboration: Moneti et al. 2022). A source catalogue is available; however, the aperture sizes are inappropriate for the shape of the BCG. Furthermore, the BCG complex is blended with many neighbouring galaxies at and and is barely detected at and . We thus estimated the flux directly from the mosaics (see Sect. 3.3).
The coordinates of the brightest core in the BCG complex correspond ( separation) to a point source detected at in the Spitzer Wide-area InfraRed Extragalactic Survey (SWIRE; see Lonsdale et al. 2003, 2004). Given the relative isolation of this source (the closest MIPS source is away) and the FHWM of the MIPS point-spread function (PSF; Rieke et al. 2004), we adopted the SWIRE flux here.
@@ 2.4 Keck MOSFIRE spectroscopy
The Euclid Q1 release contains NISP slitless multiobject spectra taken with the red grism (Euclid Collaboration: Copin et al. 2025; Euclid Collaboration: Jahnke et al. 2025). However, the reduction of the grism data was not optimal in Q1; most of the identified issues are being addressed for the Euclid DR1. In the Puddle cluster field of view, an average of 55.6% of the pixels of the calibrated unidimensional spectra (excluding the edges) are flagged for suspicious behaviour.
@@
Spectroscopic data were reduced with the PypeIt data reduction pipeline version 1.17.1 (Prochaska et al. 2020), following the standard calibration and extraction process. The reduction process included bias subtraction, flat-field correction, and wavelength calibration employing arc lamp frames. Sky subtraction was performed using the PypeIt algorithm, which models the sky background on a slit-by-slit basis to account for variations across the field of view. PypeIt outputs reduced and co-added 2D spectra, from which we extract 1D spectra using a constant extraction width as a function of wavelength with an object-finding algorithm. The extraction width is calculated based on a PypeIt-based optimal aperture mask maximising the S/N. We examined the 1D and 2D spectra to locate emission lines.
@@ 3 Analysis
The detection of the Puddle cluster in MaDCoWS2 and Euclid (Euclid Collaboration: Bhargava et al. 2025) indicated the presence of a sizeable overdensity. MOSFIRE spectroscopy provided spectroscopic redshifts centred at for the BCG (see Sect. 3.2) and for a galaxy at away from the BCG complex, which may be part of the cluster infalling region. In this section, we explored the candidate cluster properties with the photometric and spectroscopic data sets. When necessary, we assumed that the redshifts of the cluster and the BCG are .
@@ 3.1 The galaxy population
Figure 2 presents the colour-magnitude diagram (CMD) of the Puddle cluster. The galaxies out to away (about Mpc) from the proto-BCG are shown as green points, and those between and are in black. The red sequence colour is 2.44 at (pink line), assuming Solar metallicity. It is computed with Bagpipes (Carnall et al. 2018) with a model consisting of a single burst of star formation at followed by passive evolution. Combining this model with the Mancone et al. (2010) luminosity function for the IRAC channel, we compute a characteristic magnitude () of 21.4 in band. A formation redshift of 3 generates similar results: a red sequence at and .
The central structure is detected as four galaxies by the MER pipeline. However, given the irregular shape and the presence of substantial diffuse emission, the MER photometry is unreliable for these four sources. Thus, we removed those four detections from our CMDs and performed our own measurement of the proto-BCG flux in Sect. 3.3. The cyan dot in Fig. 2 represents the position of the central structure in the CMD.
The red sequence of our CMD appears sparsely populated compared with clusters confirmed at (e.g. Papovich et al. 2010; Andreon et al. 2014; Webb et al. 2015b; Strazzullo et al. 2016; Willis et al. 2020). We see in particular a lack of red galaxies at , which is unusual among the known high-redshift clusters: massive galaxies are more susceptible to internal quenching (e.g. Peng et al. 2010b) and thus tend to be redder than less massive ones even at (e.g. Kawinwanichakij et al. 2017).
@@ 3.1.1 Background subtraction and completeness correction
To assess the importance of the galaxy overdensity associated with the Puddle cluster and its location in the CMD diagram, we estimated the contribution of field galaxies and subtracted it from our CMD. To measure this ‘background’, we randomly selected 100 non-overlapping circular regions () in the Fornax field. We computed the average density of galaxies in these fields as a function of magnitude and colour and we used only the galaxies that are at least -detected in both and bands. For each magnitude and colour, we then subtracted this average density from the galaxy density in the central region of the Puddle cluster. The background subtraction uncertainties were determined from the standard deviation of the galaxy densities in the random fields.
@@
We also used the 100 random fields described above to determine completeness corrections. Our computation followed the same principle than the completeness correction described in the Appendix A.1. of van der Burg et al. (2020). Figure 3 presents a visual summary of the process. We fitted an exponential to the highly complete part of the distribution in each band empirically determined as and . It should be noted that since the -band completeness is the most critical one, we have been especially conservative in our selection of the regime which we consider highly complete. We then computed a completeness factor for magnitudes greater than 23 () or 23.5 () by taking the ratio of the observed galaxy count to the count predicted by the best fit. We assumed a completeness factor of one for and . The total 50% completeness, shown on Fig. 2 as a purple line, is estimated by multiplying the and -band completeness factor for a given set of and magnitudes. This multiplication relies on the assumption that there is no correlation between detections in the and bands. Since that may not be true, we limited the rest of our analysis to , a regime where the completeness correction does not fall below 34% and is entirely determined by the -band completeness (i.e. the completeness correction is entirely determined by colour). The completeness uncertainties were propagated from the uncertainties on the best fit parameters. For example, the correction for a -band completeness of 0.5 is .
@@
Figure 4 presents the impact of these corrections on the galaxy population of the Puddle cluster candidate. The bottom left panel shows the background-subtracted, completeness-corrected distribution. The 50% -band completeness limit is shown as a cyan line. To avoid large (and uncertain) completeness corrections, we limit the CMD in the bottom left panel to ( is equivalent to on the red sequence and to a 76% complete band). We limited the colour to , as the background-subtracted galaxy count is consistent with zero for redder colours. The other panels show the galaxy count before and after the application of the completeness correction as a function of (top left) and colour (bottom right). For , the excess count within of the BCG is galaxies above the field level. It increases to galaxies after the application of the completeness correction. The red sequence location is indicated on the top left and bottom panels. The count uncertainties are based on the propagation of the density subtraction and the completeness uncertainties when relevant.
@@
As noted above, there are galaxies ( after the completeness correction) in excess compared to the field level in the red sequence and very few luminous red galaxies outside of the BCG complex ( galaxies with ). The bulk of the overdensity is about 1.5 magnitudes bluer than the red sequence population, and fainter than . With Bagpipes, we computed that corresponds to a Myr old stellar population, assuming Solar metallicity and no attenuation. While this blue colour suggests widespread star formation, we did not attempt to determine the cluster specific SFR. The lack of bright red galaxies and the imprecise photometric redshifts of Euclid at , are probably the two main factors explaining the difference between our galaxy counts and the Euclid Collaboration: Bhargava et al. (2025) richness estimate.
@@ 3.2 Spectroscopy
We detected continuum emission of varying robustness in 14 of the 18 MOSFIRE slitlets. We also identified unambiguous isolated emission lines, most likely H , in three slitlets: one slitlet targeting the BCG complex, one slitlet targeting a source approximately from the MaDCoWS2 centroid, and another slitlet far from the MaDCoWS2 centroid that has two serendipitous emission lines. In addition, faint, tentative emission lines were detected in three other slitlets. Based on the one of these faint tentative line, we estimated a line detection threshold of at the central wavelengths of the -band grism. This corresponds to a star-formation rate of , assuming a standard conversion factor between star-formation rate and H line luminosity (Hao et al. 2011). At m, the wavelength of the emission line detected from the BCG complex, the atmosphere absorbs of the flux, implying a rough star-formation rate detection threshold of 8. The presence of faint traces and tentative emission lines suggests that the exposure time calculations were perhaps too optimistic. We also note that the outcome of our observation is comparable to Stanford et al. (2012) attempt to confirm IDCS J1426.5+3508, a cluster, with spectra from the Low-Resolution Imaging Spectrograph on Keck I. On 16 targets, they confirmed two members. However, the main challenge in our data reduction was the subtraction of the strong near-infrared sky lines (e.g. Webb et al. 2015b; Balogh et al. 2017; Belli et al. 2019), while Stanford et al. (2012) main limitation was the faintness of quiescent galaxies in the optical (which corresponds to rest-frame ultraviolet at ).
@@
The emission region in Fig. 5 is broad, spanning roughly Å and presents a complex structure. Such a broad range suggests three possibilities: a single H line from an AGN; at least two sources of H emission along the same line-of-sight; or H emission blended with the [N ii] doublet. We examined these possibilities by modelling the spectrum as one, two, and three Gaussians, weighting the contribution of each pixel by the inverse of its variance. We computed the reduced and Akaike information criterion (AIC, e.g. Akaike 1974; Liddle 2007) of each fit. The results are presented in Table 1. The middle and lowest panels of Fig. 5 show the one-, two-, and three-Gaussian decompositions in blue, pink, and purple, respectively. Due to the limited S/N of the spectrum, we did not attempt to subdivide the H emission into broad and narrow components.
The simplest case, the single Gaussian model, yields the highest reduced and AIC. Both measures strongly suggest that the two- or three-Gaussians models are better representations of the spectrum. The components of the three-Gaussian model are not treated independently: We fixed the amplitude ratio of the [N ii] doublet as , in agreement with theoretical (e.g. Galavís et al. 1997; Storey and Zeippen 2000; Bon et al. 2025) and observational (Dočinović et al. 2023) determinations, and forced its members to have the same width. We forced all three emission lines to have the same redshifts. We also tested a fit in which we relaxed the last requirement and allowed the H emission to lie at a different redshift than the [N ii] doublet. We found very similar redshifts () and the FWHM of each line was not significantly changed compared to the previous fit. We discussed the interpretation of the various fits in Sect. 4.2.1.
@@ 3.3 The BCG complex
The right panel of Fig. 1 presents a tricolour image (, , and ) of the BCG complex of the Puddle cluster, with the MOSFIRE slit and the custom aperture used for the photometry overlaid (see Sect. 3.3.1). This system has a complex morphology, with diffuse emission embedding at least six cores and extending in a kind of tail towards the southwest. There is a seventh core within the tail emission. The projected distance between cores is kpc, and in the case of the six central cores kpc. Along the axis of the diffuse tail, the system is kpc across.
@@ 3.3.1 Euclid and DECam photometry
To estimate the magnitude of the BCG complex, our first step was to determine a suitable aperture to measure the system photometry. To do so, we used sep.extract to compute the segmentation map in the NISP band and set our detection threshold at with a minimum area of 30 pixels. We then assigned a value of one to the pixels corresponding to the BCG complex detections and removed all other detections in our segmentation map. The resulting detection was very irregularly shaped, so we used a Gaussian kernel with 3 pixels to smooth the detection shape, retaining only the pixels with a value of 0.25 or more. We then used the resulting shape (shown as white contours in the right panel of Fig. 1) to measure the photometry in all Euclid and DECam bands.
To estimate the photometric uncertainties, we adopted a statistical approach: We found all detections in the image of interest, with a detection threshold at and a minimum area of 5 pixels. We then randomly displaced the footprint of the central merger until we found a position with no overlapping detections. We measured the sky flux in that position and repeated the process for 100 random empty sky patches. We took the standard deviation of the sky fluxes as our sky uncertainties, which we then added in quadrature to the Poisson noise of the aperture photometry.
We did not compute aperture corrections for DECam. As long as the defined aperture contains most of the flux in every band, the aperture corrections should be negligible. This is the case for all DECam bands. In 3 segmentation maps, the BCG complex, if detected, is entirely enclosed by our custom aperture.